OpenStudy (acacia):

Solve by the substitution method. y = 2x - 4 7x – 5y = 14

OpenStudy (anonymous):

substitute 2x-4 for y. 7x-5(2x-4)=14

OpenStudy (luigi0210):

What she said.

OpenStudy (jdoe0001):

what he said he said

OpenStudy (luigi0210):

xD

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Great answer ;)

OpenStudy (luigi0210):

@acacia can you finish it?

OpenStudy (acacia):

@Luigi0210 yeah give me a sec

OpenStudy (luigi0210):

Alright, let's see what you got kiddo :D

OpenStudy (luigi0210):

*when you finish that is :)

OpenStudy (acacia):

{Y=2x-4 {7x-5y=14 7x-5(2x-4)=14 solve for x 7x-10x+20=14 -3x=14-20 x=-6/-3 x=2 right??

OpenStudy (luigi0210):

Yup, x=2 :)

OpenStudy (luigi0210):

Now plug that back in and find y.

OpenStudy (acacia):

Then I plug it back in to Y=2(2)-4 and Y=0

OpenStudy (luigi0210):

Correct!

OpenStudy (luigi0210):

Great job. :)

OpenStudy (acacia):

Whoo break through! @Luigi0210

OpenStudy (luigi0210):

Hm?